OpenStudy (dustinleikam23):

On a backpacking trip through Europe, Troy spent 20 Swiss francs to stay one night in a youth hostel. If the foreign exchange rate between the U.S. dollar and the Swiss franc is 1:2.2, how much did Troy spend in U.S. dollars? $44 $9.09 $19 $20

RhondaSommer (rhondasommer):

\(\Huge\mathbb{WELCOME~2~OPENSTUDY}\)

RhondaSommer (rhondasommer):

\[\frac{1}{2.2}=\frac{x}{20}\]\[cross~multiply\]\[20*1=2.2*x\]\[can~you~simplify?\]

OpenStudy (bemoseemo):

Hello Fellow Plato User